# CAST: Cross-Attentive Spatio-Temporal feature fusion for Deepfake detection

### v1

@@ 5.5.4 Convolutional Projection Layer
To investigate the role of the $1\times 1$ convolutional projection layer beyond simple dimensionality reduction, we conducted a study on its impact within our CNN-transformer framework. The CAST model already uses EfficientNet-B0 with an output feature dimension of 1280 which is identical to the transformer encoder’s embedding size. This setup allowed us to evaluate the effect of the convolutional projection purely in terms of feature adaptation and stability, rather than dimensional compression. We bypass the $1\times 1$ convolution and directly feed the raw CNN feature maps into the cross-attention module. Similarly, we skipped the convolutional projection after Global Average Pooling (GAP) and directly used the pooled feature maps to prepare for the transformer encoder. From the result reported in Table 6, removing the convolution leads to a noticeable drop in performance, particularly on the more challenging datasets. The AUC on Celeb-DF v2 dropped from 73.97% to 68.14%, and on DFD from 91.67% to 88.49%, while FF++ (HQ) maintained relatively high accuracy at 98.60%. These findings highlight that, even when the dimensionalities are aligned, the convolutional projection serves a crucial role in refining feature representations and ensuring a stable attention mechanism.
@@
Model Variant AUC (%) FF++ (HQ) Celeb-DF DFD No Cross-Attn (CNN + Transformer) 98.12 66.18 90.10 Self-Attn (Spatial + Temporal sep.) 96.38 71.41 87.24 Cross-Attn reversed (Spatial $\rightarrow$ Temporal) 97.93 71.25 90.41 No projection layer 98.60 68.14 88.49 CAST-B0 98.98 76.98 93.31
@@ 5.5.5 Impact of CNN Backbone Architecture
To assess the effect of backbone architecture on spatial feature quality and downstream classification, we evaluated three CNN extractors: EfficientNet-B0, EfficientNet-B5, and XceptionNet. As summarized in Table 7, EfficientNet-B5 achieved the best overall AUC score of 99.49% on FF++(HQ). However, EfficientNet-B0 slightly outperformed it on the Celeb-DF and DFD dataset, achieving an AUC of 76.98% and 93.31% respectively, suggesting superior generalization under domain shift. XceptionNet, while commonly used in earlier works, yielded the lowest performance across all datasets.
To further understand these differences, we visualized the high-level feature representations from each backbone using t-SNE, as shown in Fig. 6. Each subplot presents the separability of real and fake samples for a given dataset–backbone pair. EfficientNet-B5 consistently produces well-separated clusters for both FF++ dataset (DF and F2F), with minimal overlap between real and manipulated distributions. EfficientNet-B0 shows slightly more dispersion yet maintains clear boundaries. In contrast, XceptionNet’s embeddings are noticeably entangled, especially on DFD, confirming its relatively weaker spatial encoding. These patterns highlight the importance of selecting a backbone that balances spatial discriminability with robustness across manipulation types and data domains.
@@
Table 7: Impact of CNN backbone on model performance. Comparison of model variants using AUC scores based on intra-dataset testing on FF++(HQ) and cross-dataset testing on Celeb-DF and DFD datasets. All models are variants of our proposed CACT model with different CNN backbones.
CNN Backbone AUC (%) FF++ (HQ) Celeb-DF DFD EfficientNet-B0 [30] 98.98 76.98 93.31 EfficientNet-B5 [30] 99.49 74.92 92.07 Xception [20] 97.82 66.04 89.10
@@ 6 Limitations
While the proposed CACT framework demonstrates strong performance in both intra- and cross-dataset evaluation, several limitations remain that merit attention.
First, in Table 2, we observe that although the model performs competitively in intra-manipulation settings achieving AUCs of 100.00% (DF) and 99.29% (FS), and achieves high AUC scores in majority of the cross-testing scenarios as well, its generalization across manipulation types is comparatively degraded in case of the FS sub-dataset. Notably, when trained solely on FS, CACT-B0 achieves only 59.97% and 51.11% AUC on DF and NT respectively. Similar trends are observed for CACT-B5, with 58.13% and 49.95% AUC on those subsets. These results indicate that representations learned from FS are highly specialized and exhibit limited transferability. This challenge extends to cross-testing conditions as well. When trained on DF, the CACT-B0 model’s AUC on FS drops to just 60.15%, as compared to 79.51% of WATCHER [34], 74.29% of RECCE [32] and 73.61% of ADAL [33] highlighting its difficulty in detecting face-swapping manipulations not seen during training.
Second, from Table 3, we note that although our models achieve competitive AUC values (98.98% for B0 and 99.49% for B5) on FF++ (HQ), the classification accuracy (ACC) lags slightly behind other state-of-the-art approaches. Specifically, CACT-B0 achieves 96.89% ACC, the lowest among listed methods, while CACT-B5 reaches 97.57%, still underperforming compared to MAT (97.60%). This discrepancy between AUC and ACC suggests that although the models rank predictions effectively, their decision thresholds may not be ideally calibrated under high-quality compression, possibly due to the subtle nature of discriminative features in such settings
@@ 7 Conclusion
In this study we presents CAST, a novel Cross-Attentive Spatio-Temporal feature fusion architecture tailored for robust deepfake video detection. Specifically, we introduce a cross-attention mechanism to effectively fuse spatial and temporal features extracted from a CNN-Transformer network. By explicitly modeling the interplay between spatial and temporal features, the proposed cross-attention mechanism enables temporal representations to attend selectively to manipulation-prone spatial regions such as those affected by unnatural warping, expression inconsistencies, or flickering artifacts. Hence, enhancing the network’s capacity to localize and detect subtle forgeries. Comprehensive evaluations are conducted across multiple benchmarks, including FaceForensics++ (HQ), Celeb- DF v2, and DFD, under a variety of testing protocols, including intra-dataset, cross-manipulation, and multi-source setups. CAST consistently outperformed and was comparable to SOTA methods in most settings. Specifically, under the high-quality single-manipulation scenario, CAST-B0 variant achieved 100% AUC on DF, 86.29% on F2F and 92.02% on NT, while CAST-B5 achieved 99.99% on FS, confirming the framework’s strong detection capabilities when trained on single manipulations. In intra-dataset evaluation, CAST-B5 achieved the highest AUC score of 99.49% as compared to the SOTA methods along with a comparable accuracy of 97.57%. In cross-manipulation and multi-source contexts, CAST demonstrated favorable generalization, outperforming SOTA approaches when cross-tested on DFD dataset achieving 93.31% AUC score, while maintaining competitive AUC scores when evaluated on unseen Celeb-DF videos. Ablation studies confirmed the critical role of cross-attention in enhancing generalization, with its removal leading to significant drops in cross-dataset performance. Replacing it with decoupled self-attention further underscored the importance of explicit spatial-temporal interaction. Directionality experiments showed that using temporal queries are more effective for aligning manipulation cues. Moreover, the convolutional projection layer is essential for stable attention computation and backbone comparisons reveal that CAST-B5 achieves higher intra-dataset accuracy whereas CAST-B0 shows better generalization, making it well-suited for efficient real-world deployment.

### v2

@@ 5.5.5 Convolutional Projection Layer
To investigate the role of the $1\times 1$ convolutional projection layer beyond simple dimensionality reduction, we conducted a study on its impact within our CNN-transformer framework. The CAST model already uses EfficientNet-B0 with an output feature dimension of 1280 which is identical to the transformer encoder’s embedding size. This setup allowed us to evaluate the effect of the convolutional projection purely in terms of feature adaptation and stability, rather than dimensional compression. We bypass the $1\times 1$ convolution and directly feed the raw CNN feature maps into the cross-attention module. Similarly, we skipped the convolutional projection after Global Average Pooling (GAP) and directly used the pooled feature maps to prepare for the transformer encoder. From the result reported in Table 6, removing the convolution leads to a noticeable drop in performance, particularly on the more challenging datasets. The AUC on Celeb-DF v2 dropped from 76.98% to 68.14%, from 93.31% to 88.49% on DFD and from 81.25% to 74.19% on DFDC, while FF++ (HQ) maintained relatively high accuracy at 98.60%. These findings highlight that, even when the dimensionalities are aligned, the convolutional projection serves a crucial role in refining feature representations and ensuring a stable attention mechanism.
@@ 5.5.6 Impact of CNN Backbone Architecture
To assess the effect of backbone architecture on spatial feature quality and downstream classification, we evaluated three CNN extractors: EfficientNet-B0, EfficientNet-B5, and XceptionNet. As summarized in Table 7, EfficientNet-B5 achieved the best overall AUC score of 99.49% on FF++(HQ). However, EfficientNet-B0 slightly outperformed it on the Celeb-DF, DFD and DFDC dataset, achieving an AUC of 76.98%, 93.31% and 81.25% respectively, suggesting superior generalization under domain shift. XceptionNet, while commonly used in earlier works, yielded the lowest performance across all datasets.
To further understand these differences, we visualized the high-level feature representations from each backbone using t-SNE, as shown in Fig. 6. Each subplot presents the separability of real and fake samples for a given dataset–backbone pair. EfficientNet-B5 consistently produces well-separated clusters for both FF++ dataset (DF and F2F), with minimal overlap between real and manipulated distributions. EfficientNet-B0 shows slightly more dispersion yet maintains clear boundaries. In contrast, XceptionNet’s embeddings are noticeably entangled, especially on DFD, confirming its relatively weaker spatial encoding. These patterns highlight the importance of selecting a backbone that balances spatial discriminability with performance across manipulation types and data domains.
@@
Table 7: Impact of CNN backbone on model performance. Comparison of model variants using AUC scores based on intra-dataset testing on FF++(HQ) and cross-dataset testing on Celeb-DF, DFD and DFDC datasets. All models are variants of our proposed CAST model with different CNN backbones.
CNN Backbone AUC (%) Year FF++ (HQ) Celeb-DF DFD DFDC EfficientNet-B0 [39] 98.98 76.98 93.31 81.25 2019 EfficientNet-B5 [39] 99.49 74.92 92.07 78.39 2019 Xception [8] 97.82 66.04 89.10 60.63 2017
@@ 6 Limitations
While the proposed CAST framework demonstrates strong performance in both intra- and cross-dataset evaluation, several limitations remain that merit attention.
First, in Table 2, we observe that although the model performs competitively in intra-manipulation settings achieving AUCs of 100.00% (DF) and 99.29% (FS), and achieves high AUC scores in majority of the cross-testing scenarios as well, its generalization across manipulation types is comparatively degraded in case of the FS sub-dataset. Notably, when trained solely on FS, CAST-B0 achieves only 59.97% and 51.11% AUC on DF and NT respectively. Similar trends are observed for CAST-B5, with 58.13% and 53.61% AUC on those subsets. These results indicate that representations learned from FS are highly specialized and exhibit limited transferability. This challenge extends to cross-testing conditions as well. When trained on DF, the CAST-B0 model’s AUC on FS drops to just 60.15%, as compared to 79.51% of WATCHER [44], 74.29% of RECCE [3] and 73.61% of ADAL [23], highlighting its difficulty in detecting face-swapping manipulations not seen during training. This limitation can be attributed to the characteristics of the FaceSwap manipulation in FF++, where artifacts are relatively subtle and exhibit limited temporal variation across frames [33]. Unlike DF and F2F, which manipulate every frame of the target sequence, FS modifies only a subset of overlapping frames, resulting in weaker and less consistently distributed temporal cues. Since CAST relies on temporal–spatial inconsistencies to guide cross-attention fusion, the reduced strength of temporal signals in FS limits the contribution of the temporal branch and degrades cross-manipulation generalization.
Second, from Table 3, we note that although our models achieve competitive AUC values (98.98% for B0 and 99.49% for B5) on FF++ (HQ), the classification accuracy (ACC) lags slightly behind other state-of-the-art approaches. Specifically, CAST-B0 achieves 96.89% ACC, the lowest among listed methods, while CAST-B5 reaches 97.57%, still underperforming compared to MAT (97.60%). This discrepancy between AUC and ACC suggests that although the models rank predictions effectively, their decision thresholds may not be ideally calibrated under high-quality compression, possibly due to the subtle nature of discriminative features in such settings.
A similar limitation is observed in the cross-dataset evaluation on Celeb-DF, where CAST achieves competitive yet comparatively lower performance than on datasets such as DFD. Celeb-DF consists of high-fidelity, artifact-sparse deepfake videos with explicitly reduced temporal flickering through the use of temporally correlated facial landmarks, resulting in smoother frame-to-frame transitions and fewer temporal inconsistencies [24]. Consequently, the reduced availability of temporal irregularities limits the effectiveness of detection models that rely on temporal cues. This challenge is further amplified by the significant distribution shift between FF++ and Celeb-DF.
@@ 7 Conclusion
In this study we presents CAST, a novel Cross-Attentive Spatio-Temporal feature fusion architecture tailored for robust deepfake video detection. Specifically, we introduce a cross-attention mechanism to effectively fuse spatial and temporal features extracted from a CNN-Transformer network. By explicitly modeling the interplay between spatial and temporal features, the proposed cross-attention mechanism enables temporal representations to attend selectively to manipulation-prone spatial regions such as those affected by unnatural warping, expression inconsistencies, or flickering artifacts. Hence, enhancing the network’s capacity to localize and detect subtle forgeries. Comprehensive evaluations are conducted across multiple benchmarks, including FaceForensics++ (HQ), Celeb-DF v2, DFD and DFDC under a variety of testing protocols, including intra-dataset, cross-manipulation, and multi-source setups. CAST consistently outperformed and was comparable to SOTA methods in most settings. Specifically, under the high-quality single-manipulation scenario, CAST-B0 variant achieved 100% AUC on DF, 86.29% on F2F and 92.02% on NT, while CAST-B5 achieved 99.99% on FS, confirming the framework’s strong detection capabilities when trained on single manipulations. In intra-dataset evaluation, CAST-B5 achieved the highest AUC score of 99.49% as compared to the SOTA methods along with a comparable accuracy of 97.57%. In cross-manipulation and multi-source contexts, CAST demonstrated favorable generalization, outperforming SOTA approaches when cross-tested on DFD and DFDC datasets, achieving 93.31% and 81.25% AUC scores respectively, while maintaining competitive performance when evaluated on unseen Celeb-DF videos. Ablation studies confirmed the critical role of cross-attention in enhancing generalization, with its removal leading to significant drops in cross-dataset performance. Replacing it with decoupled self-attention further underscored the importance of explicit spatial-temporal interaction. Directionality experiments showed that using temporal queries are more effective for aligning manipulation cues. Moreover, the convolutional projection layer is essential for stable attention computation and backbone comparisons reveal that CAST-B5 achieves higher intra-dataset accuracy whereas CAST-B0 shows better generalization, making it well-suited for efficient real-world deployment.